OpenStudy (anonymous):

HELP PLEASEEETara creates a budget for her weekly expenses. The graph shows how much money is in the account at different times. Find the slope of the line.

OpenStudy (anonymous):

heres the pic that goes with it

OpenStudy (anonymous):

i know what slope is im just convulsed on how to solve it.

OpenStudy (mayankdevnani):

Just go through it , hope it help :- 1) http://www.mathopenref.com/coordslope.html 2) http://www.mathwarehouse.com/algebra/linear_equation/slope-of-a-line.php

OpenStudy (mayankdevnani):

@luvalpaca

OpenStudy (anonymous):

it doesnt, bruh ive been to those sites ty thou

OpenStudy (anonymous):

with this it would be rise over run rise/run i believe. it would easier if u had it on graph paper

OpenStudy (anonymous):

theres a pic thou

OpenStudy (mayankdevnani):

i can help you

OpenStudy (anonymous):

i know but if it was on graph paper it would be so much easier

OpenStudy (anonymous):

wait is it 2/4

OpenStudy (mayankdevnani):

\[\huge \bf Slope=\frac{rise}{run}\]

OpenStudy (mayankdevnani):

so, \[\huge \bf Slope=\frac{y_1-y_2}{x_1-x_2}\] where x1=4,x2=12 , y1=2400 and y2=2000

OpenStudy (mayankdevnani):

finally, \[\huge \bf Slope=\frac{2400-2000}{4-12}=\frac{400}{-8}=\frac{-50}{1}\]

OpenStudy (mayankdevnani):

hope you got it . @luvalpaca

OpenStudy (anonymous):

ok thank u

OpenStudy (mayankdevnani):

welcome :)